OpenStudy (anonymous):

Paul needs 2 1/4 yards of fabric to make a tablecloth. How many tablecloths can he make from 18 1/4 yards of fabric? A. 72 tablecloths B. 10 tablecloths C. 9 tablecloths D. 8 tablecloths

OpenStudy (anonymous):

18 .25 - 2.25= 16 16-2.25=13.75-2.25=11.5-2.25= 9.25-2.25=7-2.25=4.5-2.25= 2.5-2.25=.25 so you can make 8 tablecloths, or D. glad to help!!!